# Minimum-Action Learning: Energy-Constrained Symbolic Model Selection for Physical Law Identification from Noisy Data

### 16 Mar 2026 v1

@@ Abstract
Identifying physical laws from noisy observational data is a central challenge in scientific machine learning. We present Minimum-Action Learning (MAL), a framework that selects symbolic force laws from a pre-specified basis library by minimizing a Triple-Action functional combining trajectory reconstruction ($I_{\mathrm{max}}$), architectural sparsity ($E_{\mathrm{min}}$), and energy-conservation enforcement ($\mathcal{L}_{\mathrm{Symmetry}}$). A wide-stencil acceleration-matching technique reduces noise variance by $10^{4}\times$, transforming an intractable problem (SNR $\approx 0.02$) into a learnable one (SNR $\approx 1.6$); this preprocessing is the critical enabler shared by all methods tested, including SINDy variants. Using two benchmarks—Kepler (inverse-square gravity) and Hooke’s law (linear restoring force)—we show that MAL recovers the correct force law with Kepler exponent $p=3.01\pm 0.01$ in 835 s at $\sim$0.07 kWh (40% reduction vs. prediction-error-only baselines). The raw correct-basis rate is 40% for Kepler and 90% for Hooke; an energy-conservation-based model selection criterion discriminates the true force law from alternatives in all cases, yielding 100% pipeline-level identification. A physics-informed gate initialization independently achieves 10/10 correct selection. Direct comparison against noise-robust SINDy variants, HNNs, and LNNs confirms MAL’s distinct niche: energy-constrained, interpretable model selection combining symbolic basis identification with dynamical rollout validation. The sparse architectures that emerge exhibit intrinsic crystallization timescales ($\Delta t_{\text{span}}=36.2\pm 4.1$ epochs) independent of initialization. The regularization design draws structural inspiration from biological metabolic constraints, though the biological analogies discussed remain suggestive rather than formally established.
Keywords: neural architecture search, action principles, energy optimization, physical law identification, modularity, developmental neuroscience
The principle of least action, formulated by Maupertuis and refined by Lagrange and Hamilton, has unified physics from classical mechanics to quantum field theory [14]. Simultaneously, biological systems—constrained by thermodynamics and natural selection—have evolved architectures that minimize metabolic expenditure while maximizing information processing [44, 25]. We propose that these parallel optimization strategies are not coincidental but reflect a deeper organizing principle operating across scales: network-weighted action minimization. Here we demonstrate that embedding this principle in artificial neural networks—via a Triple-Action functional combining information maximization ($I_{\mathrm{max}}$), energy minimization ($E_{\mathrm{min}}$), and symmetry enforcement—enables energy-constrained identification of physical laws from noisy data, while spontaneously generating the modular architectures characteristic of evolved systems.
Introduction
The Energy-Physics Connection
From Claude Bernard’s milieu intérieur [2] through Cannon’s homeostasis [5] to contemporary network physiology [1], biology has recognized that living systems operate through energy-constrained multi-scale integration. Kleiber’s discovery [25] that metabolic rate scales as $M^{3/4}$ across 17 orders of magnitude in body mass, later explained by West, Brown, and Enquist [44] through network optimization, demonstrates that metabolic constraints impose mathematical regularities on biological organization.
In parallel, physics has long relied on action principles—the requirement that systems follow paths extremizing $S=\int L\,dt$, where $L$ is the Lagrangian—to derive equations of motion [17]. From classical mechanics to quantum field theory, nature appears to solve optimization problems, minimizing action subject to symmetry constraints [43].
Recent work in developmental neuroscience reveals that these principles converge in the brain: glial cells enforce metabolic constraints on neural architecture search during development [19, 42], while dynamic coordination theory [34, 22, 23, 24] formalizes the relationship between energy states and network synchronization through metastable dynamics [38] and Farey-tree patterns—a framework supported by nonlinear frequency-locking phenomena in autonomic neural circuits [16] and phase-oscillator models of synaptic plasticity [6]. We hypothesize that this metabolic architecture search provides a template for energy-efficient artificial intelligence capable of discovering physical laws.
The Kepler Benchmark: Rediscovering $F=Gm_{1}m_{2}/r^{2}$
Newton’s derivation of the inverse-square law from Kepler’s observations represents one of science’s foundational achievements. Can a machine, given only noisy positional measurements of orbiting bodies, autonomously rediscover this law? This problem—which we term the ”Kepler-with-noise” benchmark—tests whether AI systems possess genuine physical intuition or merely pattern-match existing formulations.
@@ Introduction
Existing approaches address aspects of this challenge. Symbolic regression methods [40, 11] and GNN-based symbolic distillation [11, 26] can recover force laws from clean or moderately noisy data, but degrade at high noise levels. Hamiltonian Neural Networks (HNNs) [18] and Lagrangian Neural Networks (LNNs) [10] embed conservation structure but learn black-box energy functions rather than selecting interpretable symbolic forms. Physics-informed neural networks [32] require pre-specifying the equations to solve. Standalone mathematical LLMs lack inductive biases for inverse physical reasoning (SI Appendix, Section S1), though LLM-guided symbolic search pipelines offer a complementary approach. Noether’s Razor [41] and Noether’s Learning Dynamics [37] provide complementary theoretical perspectives on symmetry-driven model selection.
We present Minimum-Action Learning (MAL), a framework that embeds biological energy-optimization principles directly into neural architecture search. By training networks to minimize a Triple-Action functional combining information content ($I_{\mathrm{max}}$), metabolic cost ($E_{\mathrm{min}}$), and symmetry constraints ($\mathcal{L}_{\mathrm{Symmetry}}$), MAL performs energy-constrained model selection over a pre-specified basis library, identifying physical laws while maintaining biological-scale energy budgets. The key contribution is not open-ended symbolic discovery but rather the demonstration that explicit metabolic constraints—inspired by the brain’s glial-neural architecture—improve both the efficiency and interpretability of physical law identification from noisy data.
Results
Energy-Constrained Identification of Newton’s Law
@@ Results
(3) Bimodal Glial-Neural Optimization (BGNO). Drawing a structural analogy from the brain’s 1:1 glia-to-neuron ratio [9] and evidence that glial cells mediate metabolic constraints during neural development [7, 12], we implement a two-phase training schedule in which the Triple-Action functional drives the architecture $\mathbf{A}(t)$ along a soft-to-discrete manifold, where energy-driven sharpening ($E_{\mathrm{min}}$) encourages the emergence of discrete structural motifs from initially uniform gate distributions. During warmup (50 epochs, low regularization $\alpha_{E}=0.01$, uniform gates $\tau=1.0$), gates $A_{i}=\mathrm{softmax}(\ell_{i}/\tau)$ remain in a soft, exploratory state; during sparsification (150 epochs, ramping $\alpha_{E}\to 1.0$, annealing $\tau\to 0.05$), the $E_{\mathrm{min}}$ subsystem penalizes non-discrete architectural states, driving gate sharpening toward one-hot selection. This mechanism is analogous to temperature-annealed differentiable architecture search [45], here motivated by biological metabolic constraints rather than computational efficiency alone. We note that the glial-neural analogy is structural, not mechanistic: the regularization weight $\alpha_{E}$ does not perform functions equivalent to glial metabolic support (e.g., lactate shuttling, potassium buffering), but serves a functionally analogous role of constraining the “neural” gradient dynamics.
@@
Figure 2: Soft-to-discrete architectural crystallization. (A) Evolution of gate activation probabilities $A_{i}$ from uniform (epoch 1, soft state) to one-hot selection of the $r^{-2}$ basis (epoch 200, discrete state). This transition occurs on a soft-to-discrete architecture manifold where A_logits are sharpened via softmax temperature decay ($\tau:1\to 0.05$), driven by the $E_{\mathrm{min}}$ subsystem which penalizes non-discrete states through gate entropy $\mathcal{L}_{\mathrm{arch}}=-\sum_{i}A_{i}\log A_{i}$. (B) The two-phase BGNO regularization schedule ($\alpha_{E}$ shown in inset) separates physics identification (warmup) from architectural sparsification. Shown for a representative seed (seed 0); variability across 10 seeds is reported in SI Table S1.
@@
  • Correct basis selection: Gate probability $A_{0}=1.000$ for the $r^{-2}$ basis in the primary run (Fig. 2). Across 10 seeds, 4 of 10 directly select $r^{-2}$; the remaining seeds select $r^{-3}$ (3 seeds) or $r^{-1}$ (3 seeds), with a Noetherian diagnostic correctly discriminating the true physics in all cases (see Robustness section).
@@
  • Energy efficiency: Total training time 835 seconds, energy consumption $\sim$0.07 kWh (full system: 200 W GPU ++ $\sim$100 W CPU/RAM/cooling), representing 40% reduction vs. prediction-error-only baselines (SI Appendix, Table S2).
@@
The Noise Bottleneck and Wide-Stencil Solution
A critical insight emerged from failure analysis: naive finite-difference acceleration estimates from noisy positions have signal-to-noise ratio (SNR) $\ll 1$, rendering learning impossible. For observation noise $\sigma_{\text{pos}}$ and timestep $\Delta t$, second-derivative variance scales as $\sigma_{a}^{2}\propto\sigma_{\text{pos}}^{2}/\Delta t^{4}$. At typical scales ($\sigma_{\text{pos}}=0.016$ AU, $\Delta t=0.05$), this yields $\sigma_{a}\approx 15.7$, while true gravitational acceleration at $r=2$ AU is $\approx 0.25$—an SNR of $0.016$ (Table 1, SI Appendix).
We solved this through wide-stencil acceleration matching: computing second derivatives with stride $s=10$ reduces noise variance by $s^{4}=10{,}000\times$ while introducing only $O(s^{2}\Delta t^{2})$ systematic error ($0.3\%$ for Keplerian orbits). This transforms an intractable problem (SNR $\approx 0.02$) into a learnable one (SNR $\approx 1.6$), enabling gradient-based discovery of the force law (Methods, SI Appendix Section S2).
@@
Robustness, Success Rate, and Intrinsic Timescales
To assess reliability, we trained 10 independent models with different random seeds on identical data. All 10 successfully crystallized to a single dominant basis, though seeds 7 and 9 required extended training beyond 200 epochs (SI Appendix, Fig. S1–S4, Table S1).
Correct-basis success rate. Of 10 seeds, 4 directly selected the correct $r^{-2}$ basis (seeds 0, 1, 2, 8); 3 selected $r^{-3}$ (seeds 3, 4, 5) and 3 selected $r^{-1}$ (seeds 6, 7, 9). The raw correct-basis rate is thus 40%. However, all 10 models recovered Kepler exponent $p\in[2.995,3.006]$—a consequence of teacher-forced trajectory matching, which preserves orbital mechanics over the limited radial range tested even when the selected basis is incorrect. This means $p\approx 3.0$ alone is insufficient to discriminate the true force law; the Noetherian criterion—already intrinsic to MAL’s trifunctional via $\mathcal{L}_{\mathrm{Symmetry}}$—must be applied as a model selection diagnostic across the seed ensemble (see below). Because the symmetry term enforces energy conservation during training, models that crystallize to the correct basis exhibit superior long-horizon Hamiltonian conservation, providing a built-in discriminant. Applying this criterion across seeds achieves 100% identification of the correct physics (SI Table S2).
Basis selection interventions. To diagnose the 40% direct selection rate, we tested three interventions across 10 seeds each (SI Table S4): (i) extended warmup (100 epochs, 4/10 correct), (ii) reduced noise ($\sigma=0.005$, 4/10 correct), and (iii) physics-informed gate initialization ($\alpha_{\text{logits}}=[1.5,0,0,0,0]$, giving $r^{-2}$ approximately 50% initial gate weight). Neither longer exploration nor cleaner data improved the success rate, but biased initialization achieved 10/10 correct selection, confirming that gate competition during the warmup phase—not noise level or exploration time—is the bottleneck. This suggests a practical recommendation: when prior knowledge favors a particular basis, encoding it as a logit bias eliminates the need for cross-seed Noetherian model selection.
@@
These timescales are intrinsic to the competition dynamics between gates driven by temperature annealing and logit gradient flow, not artifacts of the specific schedule (SI Appendix, Section S3, Fig. S5).
Generalization to Hooke’s Law
To test whether MAL generalizes beyond inverse-square gravity, we applied it to Hooke’s law (F=krF=-kr, linear restoring force) using the same basis library $\{r^{-2},r^{-1},r,1,r^{-3}\}$, identical architecture, and identical training protocol (SI Appendix, Section S4A, SI Table S3). Of 10 seeds, 9 directly selected the correct $r$ basis (90% vs. 40% for Kepler), recovering spring constant $\hat{k}=0.980\pm 0.001$ (true: $k=1.0$, 2% error). The single outlier (seed 0) selected $r^{-3}$ but was correctly rejected by the Noetherian diagnostic: across all 10 seeds, the Noetherian criterion rejected gravity-family potentials ($r^{-2}$, $r^{-1}$, $r^{-3}$) in favor of the correct linear family by a factor $3\times$ in energy conservation variance.
@@
Energy-Conservation-Based Model Selection
Because MAL’s trifunctional includes an energy-conservation term $\mathcal{L}_{\mathrm{Symmetry}}=\mathrm{Var}[E]$ (inspired by Noether’s theorem linking time-translation symmetry to energy conservation), models that crystallize to the correct basis are trained to conserve energy—providing a built-in discriminant that can be applied as a model selection criterion across seeds without any additional computation. For long-horizon rollouts (5 orbital periods), we computed Hamiltonian variance $\sigma_{H}^{2}=\langle(H-\langle H\rangle)^{2}\rangle$: models selecting the correct $r^{-2}$ basis conserve energy $3\times$ better than $r^{-1}$ and $6\times$ better than $r^{-3}$ models, despite all achieving comparable short-term trajectory accuracy (SI Table S2, Fig. 4).

### 25 Mar 2026 v2

# Minimum-action learning: Energy-constrained symbolic model selection for identifying physical laws from noisy data

@@ Abstract
Identifying physical laws from noisy observational data is a central challenge in scientific machine learning. We present Minimum-Action Learning (MAL), a differentiable framework that selects symbolic force laws from a pre-specified basis library by minimizing a triple-action functional combining trajectory reconstruction, architectural sparsity, and energy-conservation enforcement. A wide-stencil acceleration-matching technique reduces noise variance by a factor of $10^{4}$, transforming an intractable problem (signal-to-noise ratio $\approx 0.02$) into a learnable one ($\approx 1.6$); this preprocessing is the critical enabler shared by all methods tested, including SINDy variants. On two benchmarks—Kepler inverse-square gravity and Hooke’s linear restoring force—MAL recovers the correct force law with Kepler exponent $p=3.01\pm 0.01$ in $835$ s at $\sim\!0.07$ kWh, a 40% reduction relative to prediction-error-only baselines. The raw correct-basis rate is 40% for Kepler (4/10 seeds) and 90% for Hooke (9/10); an energy-conservation-based model selection criterion discriminates the true force law in all cases, yielding 100% pipeline-level identification. Direct comparison against SINDy variants, Hamiltonian neural networks, and Lagrangian neural networks confirms MAL’s distinct niche: energy-constrained, interpretable model selection that combines symbolic basis identification with dynamical rollout validation within a single differentiable framework. The sparse gate architectures that emerge exhibit intrinsic crystallization timescales ($\Delta t_{\text{span}}=36.2\pm 4.1$ epochs) independent of initialization, with geometric growth rate $\gamma=1.137\pm 0.013$ per epoch.
I Introduction
The principle of least action, formulated by Maupertuis and refined by Lagrange and Hamilton, has unified physics from classical mechanics to quantum field theory [1]. A natural question is whether this principle can be turned inward: can action minimization itself serve as a regularizer for machine-learning systems tasked with identifying physical laws from data?
@@ I Introduction
Identifying force laws from noisy observational data—the inverse problem of mechanics—remains challenging because noise amplifies catastrophically under numerical differentiation, rendering standard approaches unreliable [2, 3]. Symbolic regression methods [4, 5] and GNN-based symbolic distillation [5, 6] recover force laws from clean or moderately noisy data but degrade at high noise levels. Hamiltonian neural networks (HNNs) [7] and Lagrangian neural networks (LNNs) [8] embed conservation structure but learn black-box energy functions rather than selecting interpretable symbolic forms. Physics-informed neural networks [2] require pre-specifying the governing equations. Noether’s Razor [9] and Noether’s Learning Dynamics [10] provide complementary theoretical perspectives on symmetry-driven model selection but do not address the noise bottleneck directly.
We present Minimum-Action Learning (MAL), a differentiable framework that selects symbolic force laws from a pre-specified basis library by minimizing a triple-action functional combining trajectory reconstruction ($I_{\mathrm{max}}$), architectural sparsity ($E_{\mathrm{min}}$), and energy-conservation enforcement ($\mathcal{L}_{\mathrm{Symmetry}}$). Three design elements are central: (i) a wide-stencil acceleration-matching technique that reduces noise variance by $10^{4}\times$, transforming an intractable signal-to-noise ratio (SNR $\approx 0.02$) into a learnable one (SNR $\approx 1.6$); (ii) temperature-annealed gate competition that drives a soft-to-discrete architectural transition, selecting a single basis function from the library; and (iii) an energy-conservation-based model selection criterion, grounded in Noether’s theorem, that discriminates the true force law from alternatives across independent training runs. We validate MAL on two benchmarks—Kepler inverse-square gravity and Hooke’s linear restoring force—and compare directly against SINDy variants, HNNs, and LNNs.
The regularization design draws structural inspiration from biological metabolic constraints on neural architecture [11, 12]: the sparsity-inducing $E_{\mathrm{min}}$ term plays a role analogous to wiring-cost minimization in evolved networks. While these biological analogies remain suggestive rather than formally established, the resulting computational framework stands on its own as a physics-informed method for interpretable model selection.
II Results
II.1 Energy-constrained identification of Newton’s law
@@
(3) Two-phase training schedule. We implement a schedule in which the triple-action functional drives the architecture $\mathbf{A}(t)$ along a soft-to-discrete manifold, where energy-driven sharpening ($E_{\mathrm{min}}$) encourages the emergence of discrete structural motifs from initially uniform gate distributions. During warmup (50 epochs, low regularization $\alpha_{E}=0.01$, uniform gates $\tau=1.0$), gates $A_{i}=\mathrm{softmax}(\ell_{i}/\tau)$ remain in a soft, exploratory state; during sparsification (150 epochs, ramping $\alpha_{E}\to 1.0$, annealing $\tau\to 0.05$), the $E_{\mathrm{min}}$ subsystem penalizes non-discrete architectural states, driving gate sharpening toward one-hot selection. This mechanism is closely related to temperature-annealed differentiable architecture search [13], with the additional constraint that the sparsity pressure is explicitly tied to an energy-minimization objective inspired by wiring-cost minimization in biological networks [11, 12].
@@
Figure 2: Soft-to-discrete architectural crystallization. (A) Evolution of gate activation probabilities $A_{i}$ from uniform (epoch 1, soft state) to one-hot selection of the $r^{-2}$ basis (epoch 200, discrete state). This transition occurs on a soft-to-discrete architecture manifold where A_logits are sharpened via softmax temperature decay ($\tau:1\to 0.05$), driven by the $E_{\mathrm{min}}$ subsystem which penalizes non-discrete states through gate entropy $\mathcal{L}_{\mathrm{arch}}=-\sum_{i}A_{i}\log A_{i}$. (B) The two-phase regularization schedule ($\alpha_{E}$ shown in inset) separates physics identification (warmup) from architectural sparsification. Shown for a representative seed (seed 0); variability across 10 seeds is reported in Supplemental Table S1.
@@
  • Correct basis selection: Gate probability $A_{0}=1.000$ for the $r^{-2}$ basis in the primary run (Fig. 2). Across 10 seeds, 4 of 10 directly select $r^{-2}$; the remaining seeds select $r^{-3}$ (3 seeds) or $r^{-1}$ (3 seeds), with a energy-conservation diagnostic correctly discriminating the true physics in all cases (see Robustness section).
@@
  • Energy efficiency: Total training time 835 seconds, energy consumption $\sim$0.07 kWh (full system: 200 W GPU ++ $\sim$100 W CPU/RAM/cooling), representing 40% reduction vs. prediction-error-only baselines (Supplemental Material, Table S2).
@@
II.2 The noise bottleneck and wide-stencil solution
A critical insight emerged from failure analysis: naive finite-difference acceleration estimates from noisy positions have signal-to-noise ratio (SNR) $\ll 1$, rendering learning impossible. For observation noise $\sigma_{\text{pos}}$ and timestep $\Delta t$, second-derivative variance scales as $\sigma_{a}^{2}\propto\sigma_{\text{pos}}^{2}/\Delta t^{4}$. At typical scales ($\sigma_{\text{pos}}=0.016$ AU, $\Delta t=0.05$), this yields $\sigma_{a}\approx 15.7$, while true gravitational acceleration at $r=2$ AU is $\approx 0.25$—an SNR of $0.016$ (Table 1, Supplemental Material).
We solved this through wide-stencil acceleration matching: computing second derivatives with stride $s=10$ reduces noise variance by $s^{4}=10{,}000\times$ while introducing only $O(s^{2}\Delta t^{2})$ systematic error ($0.3\%$ for Keplerian orbits). This transforms an intractable problem (SNR $\approx 0.02$) into a learnable one (SNR $\approx 1.6$), enabling gradient-based discovery of the force law (Methods, Supplemental Material Section S2).
@@ II.2 The noise bottleneck and wide-stencil solution
II.3 Robustness, success rate, and intrinsic timescales
To assess reliability, we trained 10 independent models with different random seeds on identical data. All 10 successfully crystallized to a single dominant basis, though seeds 7 and 9 required extended training beyond 200 epochs (Supplemental Material, Fig. S1–S4, Table S1).
Correct-basis success rate. Of 10 seeds, 4 directly selected the correct $r^{-2}$ basis (seeds 0, 1, 2, 8); 3 selected $r^{-3}$ (seeds 3, 4, 5) and 3 selected $r^{-1}$ (seeds 6, 7, 9). The raw correct-basis rate is thus 40%. However, all 10 models recovered Kepler exponent $p\in[2.995,3.006]$—a consequence of teacher-forced trajectory matching, which preserves orbital mechanics over the limited radial range tested even when the selected basis is incorrect. This means $p\approx 3.0$ alone is insufficient to discriminate the true force law; the energy-conservation criterion—already intrinsic to MAL’s trifunctional via $\mathcal{L}_{\mathrm{Symmetry}}$—must be applied as a model selection diagnostic across the seed ensemble (see below). Because the symmetry term enforces energy conservation during training, models that crystallize to the correct basis exhibit superior long-horizon Hamiltonian conservation, providing a built-in discriminant. Applying this criterion across seeds achieves 100% identification of the correct physics (Supplemental Table S2).
Basis selection interventions. To diagnose the 40% direct selection rate, we tested three interventions across 10 seeds each (Supplemental Table S4): (i) extended warmup (100 epochs, 4/10 correct), (ii) reduced noise ($\sigma=0.005$, 4/10 correct), and (iii) physics-informed gate initialization ($\alpha_{\text{logits}}=[1.5,0,0,0,0]$, giving $r^{-2}$ approximately 50% initial gate weight). Neither longer exploration nor cleaner data improved the success rate, but biased initialization achieved 10/10 correct selection, confirming that gate competition during the warmup phase—not noise level or exploration time—is the bottleneck. This suggests a practical recommendation: when prior knowledge favors a particular basis, encoding it as a logit bias eliminates the need for cross-seed energy-conservation model selection.
@@ II.3 Robustness, success rate, and intrinsic timescales
These timescales are intrinsic to the competition dynamics between gates driven by temperature annealing and logit gradient flow, not artifacts of the specific schedule (Supplemental Material, Section S3, Fig. S5).
II.4 Generalization to Hooke’s law
To test whether MAL generalizes beyond inverse-square gravity, we applied it to Hooke’s law (F=krF=-kr, linear restoring force) using the same basis library $\{r^{-2},r^{-1},r,1,r^{-3}\}$, identical architecture, and identical training protocol (Supplemental Material, Section S4A, Supplemental Table S3). Of 10 seeds, 9 directly selected the correct $r$ basis (90% vs. 40% for Kepler), recovering spring constant $\hat{k}=0.980\pm 0.001$ (true: $k=1.0$, 2% error). The single outlier (seed 0) selected $r^{-3}$ but was correctly rejected by the energy-conservation diagnostic: across all 10 seeds, the energy-conservation criterion rejected gravity-family potentials ($r^{-2}$, $r^{-1}$, $r^{-3}$) in favor of the correct linear family by a factor $3\times$ in energy conservation variance.
@@ II.4 Generalization to Hooke’s law
II.5 Energy-conservation-based model selection
Because MAL’s trifunctional includes an energy-conservation term $\mathcal{L}_{\mathrm{Symmetry}}=\mathrm{Var}[E]$ (inspired by Noether’s theorem linking time-translation symmetry to energy conservation), models that crystallize to the correct basis are trained to conserve energy—providing a built-in discriminant that can be applied as a model selection criterion across seeds without any additional computation. For long-horizon rollouts (5 orbital periods), we computed Hamiltonian variance $\sigma_{H}^{2}=\langle(H-\langle H\rangle)^{2}\rangle$: models selecting the correct $r^{-2}$ basis conserve energy $3\times$ better than $r^{-1}$ and $6\times$ better than $r^{-3}$ models, despite all achieving comparable short-term trajectory accuracy (Supplemental Table S2, Fig. 4).